Consider the perfectly competitive market for diesel . The aggregate demand for gasoline is the demand for gasoline

Qd=100-p while the aggregate supply is Qs =3p work out the equilibrium price and quantity


1
Expert's answer
2021-09-07T07:39:02-0400

Equilibrium price:

Prot maximising level of output and price in perfect competition is where the demand and supply are equal.


"Qd=Qs"

"100-p=3p"

"100=4p"

"p=25"


Equilibrium Quantity :

Plugging in the value of price into the demand function we get,

"Qd=100-p"

"Qd=100-25"

"Qd=75" units =supply